In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive.
(a) Construct a 95 percent confidence interval for the population proportion of positive drug tests.
(b) Why is the normality assumption not a problem, despite the very small value of p?

Answer by stanbon(57214)

In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive.
(a) Construct a 95 percent confidence interval for the population proportion of positive drug tests.
sample proportion = 1143/86991 = 0.01314
standard error = 1.96*sqrt[0.01314*0.98696/86991] = 0.00038609
---
95% CI: 0.01314-0.00038609 < p < 0.01314+0.00038509
------------------------
(b) Why is the normality assumption not a problem, despite the very small value of p?
Ans: n is large enough
============================
Cheers,
Stan H.